A tsunami causes the ocean water at a place to first go down from its normal depth of 9 m to its lowest depth of 1.5 m in a time of 4 minutes. It then rises to its maximum depth of 15.5 m before receding to its normal depth, and thereafter repeating this periodic water movement.
1
Expert's answer
2020-05-04T18:35:27-0400

A tsunami causes the ocean water at a place to first go down from its normal depth of 9 m to its lowest depth of 1.5 m in a time of 4 minutes.

"v=\frac{9-1.5}{4}=\frac{7.5}{4}=1.875m/min"

It then rises to its maximum depth of 15.5 m before receding to its normal depth, and thereafter repeating this periodic water movement

"Displacement=0"

Distance travelled in downward movement is 7.5 m(9m-1.5m).

Distance travelled in upward movement is 6.5 m(15.5m-9m).

"Distance =2*7.5+2*6.5=13+15=28 m"